OpenStudy (anonymous):

Hello!Did anybody experience difficulties in logging in?I had to make a new account just to ask a question,thus the username. Anyway,here's my question: Pls help me integrate this. {Integral Sign} (ln^4 sin [x])/(tan [x]) Please help me.

OpenStudy (anonymous):

you are not allowed to make a new account

OpenStudy (anonymous):

can you use equation editor?

OpenStudy (anonymous):

u = ln sin x du = 1/sin x * cos x = 1 / tan x

OpenStudy (anonymous):

so integral u^4 du = u^5/5 + c

OpenStudy (anonymous):

(ln (sin (x))^5 / 5 + C

OpenStudy (anonymous):

give me medal

OpenStudy (anonymous):

@barry2: I DID. DEAL WITH IT. I'M IN DESPERATE NEED HERE. Unless you can figure out how to salvage my first account, stop being so uptight about it? All my frickin medals, gone! You have no idea how I hated not opening my old account.

OpenStudy (anonymous):

i gave you the solution, where is my medal

OpenStudy (anonymous):

i gave you explanation, dont give me that lip

OpenStudy (anonymous):

@barry: Thank you. I'm very grateful for any helpful answer. And as for your medal, I'll absorb this first before I do that. Sheez, stop being so bossy. I have to understand your solution before I could actually conclude that it was a good answer.

OpenStudy (anonymous):

im not going to help you . our business is concluded until a medal is received .

OpenStudy (anonymous):

as a token of my good nature i have graciously bestowed a medal on you

OpenStudy (anonymous):

i feel like i got ripped off